OpenStudy (anonymous):

Calculate the equation of the line that is parallel to the line y = 4x - 6 and passes through the point ( 5, 8 ).

OpenStudy (anonymous):

The line parrallel to y=4x=6 has a slope of -1/4, since a parrelll line's slope is the negative reciprocal of the other line. Now that you have the slope, you can use the point slope formula y-y1=m(x-x1)

OpenStudy (anonymous):

y-8=-1/4(x-5)

OpenStudy (anonymous):

y=-1/4x+37/4